OpenStudy (anonymous):

How many ways are there to rearrange the letters in the word aptitude, if the first and last letter must each be a vowel?

OpenStudy (anonymous):

I know I have to calculate how many letters to choose from. then multiply them by each other for each spot right?

OpenStudy (anonymous):

5*26^6**5

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

lets make it easier and say the first two letters have to be vowels (clearly makes no difference) then4 choices for the first 3 for the second then you have 6 letters left the number of ways to arrange 6 letters if they are all different is \(6!=6\times 5\times 4\times 3\times 2\) but since you cannot tell the two "t" s apart, it is \(\frac{6!}{2}=6\times 5\times 4\times 3\)

OpenStudy (anonymous):

yuuup I was wrong

OpenStudy (anonymous):

multiply all this mess together to get your answer

OpenStudy (anonymous):

doing so now

OpenStudy (anonymous):

450? but I don't think that's right. working in my head not good.

OpenStudy (anonymous):

360.

OpenStudy (zarkon):

you have yet to multiply all the numbers satellite73 gave to you

OpenStudy (anonymous):

yup im lost

OpenStudy (zarkon):

4*3*6*5*4*3

OpenStudy (anonymous):

4320